reasoning · 2026-09-13
You flip a fair coin exactly three times, independently. What is the probability that all three flips are heads?
Answer: b
The chance of a head on one flip is 1/2. For three independent flips, you multiply the probabilities: 1/2 × 1/2 × 1/2 = 1/8.
Model answers
- gpt-oss-20b (openai/gpt-oss-20b): b
- qwen-3-8-27b (qwen/qwen3.8-27b): b
- nemotron-3-super (nvidia/nemotron-3-super-120b-a12b:free): b
- gemini-3-8-flash (gemini-3.8-flash-low): b
- gemini-3-1-pro (gemini-3.1-pro-high): b
- claude-opus-4-6 (claude-opus-4-6-thinking): b